# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v3

@@ IV The connection of QS to integrable PDEs: Painlevé property
We will now look for $\mathcal{B}$ with the PP such that there are no critical singularity in the entire complex plane since their very presence can affect $\mathcal{B}$ on the real line. We now recall some classical results on ODEs with the PP (details given in (Ince, 1956; Hille, 1997; Conte, 2012)). Let us rewrite the $\partial_{\ell}B$ equation in the complex plane as
@@
The only first-order ODE of the first degree that possesses the PP is the Ricatti equation, where $\mathcal{F}$ is quadratic in $\mathcal{B}$. For first-order ODEs of degree $n>1$, the equations of the form
$\mathcal{B}^{\prime}(\mathcal{Z})^{2}=P_{m}(\mathcal{B}(\mathcal{Z}))$ (50)
with $P_{m}$ a polynomial of degree $m$ possess the PP provided $n\leq 2m$. Furthermore, only four possibilities exhaust all ODEs of the form (50). With suitable variable transformations, all of them can be reduced to the following ODE with a cubic nonlinearity
@@
which is exactly solvable in terms of elliptic functions and their degenerate cases. As is well-known, elliptic functions are doubly-periodic, meromorphic functions. The dependence on the arbitrary initial condition or the TW frame thus leads to only movable poles and no movable critical singularities. Hence, the solution possesses the PP. These classical results follow from standard Painlevé test and extensions of Malmquist’s theorem (Hille, 1997; Eremenko, 1982). They show how stringent and rigid the constraint of single-valuedness can be.
@@
Returning to the QS problem, we first observe that the Riccati equation for $B$ is inconsistent with the NAE solution. The Riccati equation for $B$ will have constant real coefficients on a flux surface, which does not admit nontrivial periodic solutions on the real line Coddington et al. (1956). As discussed in Section III.1 and III.2 the perturbative treatment near the slab or the axis shows that $(\partial_{\ell}B)^{2}=P_{3}(B),P_{4}(B)$ is more appropriate, which coincides with two of the four forms of ODEs of first order and second degree with PP. The periodicity condition can be easily satisfied with the elliptic functions for these cases. Now, the cubic and quartic polynomials are the twice-integrated forms of the KdV and Gardner’s equations. The PP of the KdV and Gardner’s equation in the TW frame is well-known and is conjectured to be essential for integrable PDEs with soliton solutions (Ablowitz et al., 1980a, b).
@@
We now address whether PP is necessary and sufficient for a quasisymmetric $B$. It is certainly sufficient for QS since both the triple product form (8) and the periodicity condition can be satisfied. However, the necessity is not apparent, particularly in light of axisymmetry, where no such stringent restriction on $\partial_{\ell}B$ appears. In axisymmetry, weak singularities such as $R^{m}\log{R}^{n}$, with integer $m,n$ often arise in the description of $\psi$ (Cerfon and Freidberg, 2010) and in $B$ in cylindrical coordinates. However, these singularities are fixed and, therefore, not of the critical movable type. To further understand the difference between a perfectly axisymmetric system and a system with excellent but not exact QS, we need to look deeper into the $\bm{u}\cdot\bm{\nabla}B=0$ condition. As discussed in Appendix A, if QS were exact, we would have the same characteristics as axisymmetry. However, unlike the axisymmetric problem, the QS problem is intrinsically overdetermined and three-dimensional. With this in mind, we now motivate the PP as a criterion that allows one to look for robust quasisymmetric systems.
@@
In axisymmetry, the geometry and the $B$ decouples. Thus, all field lines are equivalent, and the TW frame, where $H=0$, is the same for all $\alpha$. Regarding the effect on the singularities, this implies that the dependence of $B$ on the field line label is precisely the same for each singularity, which amounts to a trivial rigid shift of frame. However, in a generic approximate 3D QS, the geometry and $B$ are strongly coupled. Due to the overdetermined nature, the $\bm{u}\cdot\bm{\nabla}B=0$ condition can only be approximately satisfied. As a consequence, the TW frame $H=0$ is, in general, different for different field lines. Thus, the $\alpha$ dependence of the singularities of $\partial_{\ell}B$ will vary from field line to field line. Since these movable singularities can be critical, single-valuedness of $B$ will be harder to achieve. However, when we demand the PP, the moving singularities are not critical, which respects the single-valuedness of $B$. As will be shown in Sections VI and VII, every numerically optimized quasisymmetric stellarator has the PP. Furthermore, as the optimization progresses, the PP property emerges quickly and is maintained. Therefore, we can argue that the PP is necessary and sufficient for obtaining robust QS in a 3D stellarator using numerical optimization methods, where single-valuedness and periodicity of $B$ are enforced directly using a Fourier representation.
@@
Finally, we note that none of the above arguments required knowledge of the exact force balance condition. From the perturbative calculations, we expect that force balance determines the optimal shear profile. The roots of the polynomial can vary with the flux surface label, and force balance determines their profiles.
@@
Figure 1: The value of $(dB/d\ell)^{2}$ on the outermost flux surface in the precise QA equilibrium Landreman and Paul (2022) is plotted as a function of $B$, alongside a cubic polynomial fit in (a). In (b), the errors of quadratic, cubic, and quartic fits are compared on each flux surface, demonstrating that a cubic ansatz (as predicted by KdV) is sufficient. Hence, precise QA has the Painlevé property for $B$.
@@
Figure 2: The value of $B$ in a quasiaxisymmetric device in the core ($s=1/8$) and at the edge ($s=1$). The separatrix is outside of the device but close to its edge. Note the increasing period as the separatrix is approached.
@@
Figure 3: Prediction of maximum toroidal volume that can be made quasisymmetric for precise QA. The boundary of the extended precise QA and two other flux surfaces are shown inside a Poincare plot of the islands in (a), and the roots are plotted in (b). The boundary of the original precise QA is shown in red. Note that the islands occur where the second and third roots touch approximately.
V Reduced dimensionality of quasisymmetric B
@@
In the previous Section, we have shown that $(\partial_{\ell}B)^{2}$ is a cubic or quartic polynomial in $B$, whose roots can only be functions of $\psi$. Therefore, the PP significantly reduces the dimensionality of quasisymmetric $B$, which we shall now fully explore. We shall restrict ourselves to the cubic or the KdV case since that was shown in the perturbative approach to lead to better force-balance.
With the TW constraint, the twice integrated KdV equation can be written as
@@
where the roots of the cubic polynomial, $B_{i}=\{B_{M},B_{m},B_{X}\}$, are functions of $\psi$, and $\mathcal{D}(\psi)$ is a proportionality factor. The exact solution of (52) subject to the periodicity constraint (11) can be expressed in terms of elliptic functions (Kamchatnov, 2000; Novikov et al., 1984).
To obtain the proportionality factors between $(x,t)$ and $(l,\alpha)$ explicitly, we now proceed from (52) and derive the KdV equation. Differentiating (52) twice with $\ell$ and using the $\partial_{\alpha}B$ equation (12) we get (41), the KdV equation for $B$. Here, the coordinates $(x,t)$ are defined in terms of $(\alpha,\ell)$ as
[…]
The $\partial_{\alpha}B$ relation (12), now takes the following form of a 1D traveling wave equation with speed $c(\psi)$,
@@
$B$ is then obtained from the cnoidal solution (Kamchatnov, 2000) of KdV,
$B$ $=B_{m}+(B_{M}-B_{m})\text{cn}^{2}\left(\sqrt{\frac{B_{M}-B_{X}}{2}}(x-ct),m\right),$
$m$ $=\frac{B_{M}-B_{m}}{B_{M}-B_{X}},\quad 0\leq m\leq 1.$ (55)
Here, $\text{cn}(u,m)$ is the Jacobi cosine elliptic function with modulus $m$. $B$ varies between $B_{m}$ and $B_{M}$. The period of $B$, $L(\psi)$, diverges logarithmically as $m\to 1$. One would expect this behavior near a separatrix, where periodicity along a field line is lost because the field line cannot cross the X-point. In particular, it is interesting to note that in the $m\to 1$ limit, $B$ takes the form of the infinite KdV soliton (reflectionless potential). The adiabatic invariant and the integrals (17) obtained from KdV potentials are manifestly time ($\alpha$) independent (Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990; Brezhnev, 2008).
We first note that on any flux surface $\psi$, we have four flux functions that determine $B$, namely $\mathcal{D}$ and $B_{M},B_{m},B_{X}$, the three real roots of the cubic polynomial in (52). However, as shown in Appendix D, $\mathcal{D}$ can be expressed in terms of the rotational transform $\iota$ and the roots. Therefore, the connection with KdV points to a hidden low-dimensionality of the description of $B$.
The functions $B_{M}$ and $B_{m}$ denote the maximum and minimum values of $B$ on a flux surface, respectively, and $B_{X}$ is a third root of the polynomial. The extrema of $B$, given by $B_{M}$ and $B_{m}$ depend only on $\psi$, which implies that they can be straightened out using straight field line coordinates. The behavior of $B$ in Boozer coordinates is well known (Garren and Boozer, 1991a, b; Landreman and Sengupta, 2018) near the magnetic axis and is consistent with (52). On the axis $B$ is a constant, $B_{0}$, and varies as $B\approx B_{0}(1+\overline{\eta}\sqrt{2\psi}\cos\vartheta)+2\psi(B_{20}+B_{22}\cos 2\vartheta)$ near the axis. Here, $\vartheta$ is the well-known Boozer angle (Landreman and Sengupta, 2018). The constant $\overline{\eta}$ dominates the maximum variation of $B$ on a flux surface, with the $O(2\psi)$ term being only a small second-order correction. Within the NAE framework, $B_{X}$ and $\mathcal{D}$ are constant throughout the volume, with $\mathcal{D}\propto((\iota-N/M))^{2}$. The NAE predictions are
@@
Since $B_{22},B_{0},\overline{\eta}$ are all constants related to the $B$ spectrum in Boozer coordinates, the NAE prediction (56) for the third root is a constant. However, as shown earlier in Section III.1, the third root depends on the rotational transform and, hence, the shear profile. Thus, the NAE predictions are accurate only when the rotational transform is effectively constant. The limiting surface where $m\to 1$ corresponds to $B_{X}\to B_{m}$ is the separatrix. We can use (56) to approximately predict its location.
We now look at several possibilities to understand the physical mechanisms underlying the divergence of the connection length. Broadly, these scenarios can be understood as the breakdown of the global straight field-line coordinates, the most important case being the breakdown of flux surfaces. Another possible scenario is the formation of magnetic puddles (Rodriguez and Plunk, 2023).
When the rotational transform is rational, X-points can form, near which the connection length diverges. X-points lead to the breakdown of the nestedness of flux surfaces and the formation of islands. If QS is exact, X-points can form only when $\iota$ resonates with the helicity of QS (Rodríguez and Bhattacharjee, 2021). However, even for minor QS errors, small resonant magnetic fields can lead to the formation of X-points and islands. For irrational $\iota$, sharp structures on the flux surfaces (”ridges”) can also arise (Bader et al., 2019; Punjabi and Boozer, 2020). These sharp structures can hinder magnetic field lines from crossing them (Boozer, 2015). Finally, topological obstructions such as the “Reeb components” (Duignan et al., 2024) can also lead to the non-existence of global straight field-line coordinates. However, the Reeb case requires the toroidal magnetic field to vanish at some point if MHS force balance is imposed (Duignan et al., 2024). Similarly, magnetic field strength topology breakdown can be avoided if QS does not change drastically or by numerically filtering out any non-QS part of $B$.
@@
In Section VI, we indeed observe the formation of an X-point as we approach the limiting flux surface. In practical terms, this means that QS stellarators may naturally form sharp edges as the volume of quasisymmetry is pushed to its limit, e.g., for compact devices. Sharp edges are essential for non-resonant divertors (Boozer, 2015; Punjabi and Boozer, 2020). This could have significant beneficial consequences for stellarator optimization, since these divertors would essentially be obtained for free by optimizing for quasisymmetry in compact devices.
VI Numerical verification

### v4

@@ III The connection of QS to integrable PDEs: Painlevé property
We now look for $\mathcal{B}$ with the PP such that there are no critical singularities in the entire complex plane since their very presence can affect $\mathcal{B}$ on the real line. We now recall some classical results on ODEs with the PP (details given in (Ince, 1956; Hille, 1997; Conte, 2012)). Let us rewrite the $\partial_{\ell}B$ equation in the complex plane as a first-order ODE
@@
The classification of the ODEs according to the PP depends on the degree of the ODE, or equivalently the algebraic power of the $\mathcal{B}^{\prime}(\mathcal{Z})$. The only first-order ODE of the first degree that possesses the PP is the Riccati equation, where $\mathcal{F}$ is quadratic in $\mathcal{B}$.
For first-order ODEs of degree $n>1$, the equations of the form
$\mathcal{B}^{\prime}(\mathcal{Z})^{n}=P_{m}(\mathcal{B}(\mathcal{Z}))$ (20)
with $P_{m}$ a polynomial of degree $m$, possess the PP provided $n\leq 2m$. Furthermore, only four possibilities exhaust all ODEs of the form (20). With suitable variable transformations, all of them can be reduced to the following ODE with a cubic nonlinearity
@@
which is exactly solvable in terms of elliptic functions and their degenerate cases. As is well-known, elliptic functions are doubly-periodic, meromorphic functions. The dependence on the arbitrary initial condition or the TW frame thus leads to only movable poles and no movable critical singularities. Hence, the solution possesses the PP. These classical results follow from the standard Painlevé test and extensions of Malmquist’s theorem (Hille, 1997; Eremenko, 1982). They show how stringent and rigid the constraint of single-valuedness can be.
@@
Returning to the QS problem, we first observe that the Riccati equation for $B$ is inconsistent with the NAE solution. The Riccati equation for $B$ has constant real coefficients on a flux surface, which does not admit nontrivial periodic solutions on the real line Coddington et al. (1956). For degrees $n>1$, the periodicity condition can be easily satisfied with the elliptic functions for the cubic and quartic cases. Now, the cubic and quartic polynomials are the twice-integrated forms of the KdV and Gardner’s equations. The PP of the KdV and Gardner’s equation in the TW frame is well-known and is conjectured to be essential for integrable PDEs with soliton solutions (Ablowitz et al., 1980a, b).
@@
We now address the question: is the PP necessary and sufficient for a quasisymmetric $B$? It is certainly sufficient for QS since both the triple product form (8) and the periodicity condition can be satisfied. However, the necessity is not apparent, particularly in light of axisymmetry, where no such stringent restriction on $\partial_{\ell}B$ appears. In axisymmetry, weak singularities such as $R^{m}\log{R}^{n}$, with integer $m,n$, often arise in the description of $\psi$ (Cerfon and Freidberg, 2010) and in $B$ in cylindrical coordinates. However, these singularities are fixed and, therefore, not of the critical movable type. To further understand the difference between a perfectly axisymmetric system and a system with excellent but not exact QS, we need to look deeper into the $\bm{u}\cdot\bm{\nabla}B=0$ condition. As discussed in Appendix A, if QS were exact, we would have the same characteristics as axisymmetry. However, unlike the axisymmetric problem, the QS problem is intrinsically overdetermined and three-dimensional. With this in mind, we now motivate the PP as a criterion that allows one to look for robust quasisymmetric systems.
@@
In axisymmetry, all field lines are equivalent, and the TW frame, where $H=0$, is the same for all $\alpha$. Regarding the effect on the singularities, this implies that the dependence of $B$ on the field line label is precisely the same for each singularity, which amounts to a trivial rigid shift of the frame. However, in a generic approximate 3D QS, the geometry and $B$ are strongly coupled. Due to the overdetermined nature, the $\bm{u}\cdot\bm{\nabla}B=0$ condition can only be approximately satisfied. As a consequence, the TW frame $H=0$ is, in general, different for different field lines. Thus, the $\alpha$ dependence of the singularities of $\partial_{\ell}B$ varies from field line to field line. Since these movable singularities can be critical, single-valuedness of $B$ is harder to achieve. However, when we demand the PP, the moving singularities are not critical, which respects the single-valuedness of $B$. As shown in Sections V and VI, every numerically optimized quasisymmetric stellarator has the PP. Furthermore, as the optimization progresses, the PP emerges quickly and is maintained. Therefore, we argue that the PP is necessary and sufficient for obtaining robust QS in a 3D stellarator using numerical optimization methods, where single-valuedness and periodicity of $B$ are enforced directly using a Fourier representation.
@@
Finally, we note that none of the above arguments required knowledge of the exact force balance condition. The roots of the polynomial in (21) that determine the zeros of $(\partial_{\ell}B)$ can vary with the flux surface label, and force balance determines their profiles. We explore the behavior of the roots of $(\partial_{\ell}B)^{2}$ in the next section, in the context of numerically optimized quasisymmetric stellarator equilibria.
IV Reduced dimensionality of quasisymmetric $B$
@@
Figure 2: The value of $(dB/d\ell)^{2}$ on the outermost flux surface in the precise QA equilibrium Landreman and Paul (2022) is plotted as a function of $B$, together with a cubic polynomial fit in (a). In (b), the errors of quadratic, cubic, and quartic fits are compared on each flux surface, demonstrating that a cubic ansatz (as predicted by KdV) is sufficient. Hence, precise QA has the Painlevé property for $B$.
@@
Figure 3: The value of $B$ in a quasiaxisymmetric device in the core ($s=1/8$) and at the edge ($s=1$). The separatrix is outside the device but close to its edge. Note the increasing period as the separatrix is approached.
@@
Figure 4: Prediction of the maximum toroidal volume that can be made quasisymmetric for precise QA. The boundary of the extended precise QA and two other flux surfaces are shown inside a Poincaré plot of the islands in (a), and the roots are plotted in (b). The boundary of the original precise QA is shown in red. Note that the islands occur where the second and third roots approximately touch.
@@
In the previous Section, we have shown that $(\partial_{\ell}B)^{2}$ is a cubic or quartic polynomial in $B$, whose roots can only be functions of $\psi$. Therefore, the PP significantly reduces the dimensionality of quasisymmetric $B$, which we shall now fully explore. We restrict ourselves to the cubic case,
@@
where the roots of the cubic polynomial, $B_{i}=\{B_{M},B_{m},B_{X}\}$, are functions of $\psi$, and $\mathcal{D}(\psi)$ is a proportionality factor. Since $\bm{B}\cdot\bm{\nabla}B$ vanishes at the roots, the functions $B_{M}$ and $B_{m}$ denote the maximum and minimum values of $B$ on a flux surface, while $B_{X}$ is a third root of the polynomial, possibly a saddle. The exact solution of (22), subject to the periodicity constraint (11), can be expressed in terms of elliptic functions (Kamchatnov, 2000; Novikov et al., 1984) as
$B(\ell)$ $=B_{X}+(B_{M}-B_{X})\,\mathrm{dn}^{2}\!\!\left(\sqrt{\frac{\mathcal{D}}{2}\frac{(B_{M}-B_{X})}{2}}\,\ell,\,m\right),$
$m$ $=(B_{M}-B_{m})/(B_{M}-B_{X}),\;\;0\leq m\leq 1,$ (23)
where $\mathrm{dn}(w,m)$ is the Jacobi delta-amplitude elliptic function. Note that, due to translational symmetry in $\alpha$, we can replace $\ell$ by $\overline{\ell}$ given by (14).
Equation (22) and its solution contain four functions of $\psi$: the three roots $\{B_{M},B_{m},B_{X}\}$ where $\bm{B}\cdot\bm{\nabla}B$ vanish, and the prefactor $\mathcal{D}(\psi)$. We now show that, if flux surfaces are nested and Boozer coordinates can be defined, $\mathcal{D}$ is not an independent free function but is fully determined by the roots and the rotational transform $\iota(\psi)$, reducing the description of $B$ on a flux surface to exactly three spectral parameters. We describe our reasoning below.
Since $B$ is periodic with period $L(\psi)$, integrating the field-line equation in Boozer coordinates over one full period gives a net change of $2\pi$ in the Boozer angle $\vartheta$
$\int_{0}^{L}\frac{\partial\vartheta}{\partial\ell}\,d\ell=2\pi.$ (24)
Using $\bm{B}\cdot\bm{\nabla}\vartheta=B\partial_{\ell}\vartheta$ and the Boozer Jacobian relation $\bm{B}\cdot\bm{\nabla}\vartheta=\iota_{N}B^{2}/(G+\iota I)$ (Helander, 2014), we obtain
$\oint B\,d\ell=\frac{2\pi(G+\iota I)}{\iota_{N}},$ (25)
where $\iota_{N}\equiv\iota-N$, and $G(\psi)$ and $I(\psi)$ are the poloidal and toroidal current flux functions. For purely vacuum fields $I=0$, but we keep the general expressions here.
Integrating over one period using the identity $\int_{0}^{K(m)}\mathrm{dn}^{2}(w,m)\,dw=E(m)$ yields
$\oint B\,d\ell=L\left(B_{X}+(B_{M}-B_{X})\frac{E(m)}{K(m)}\right),$ (26)
where $K(m)$ and $E(m)$ are the complete elliptic integrals of the first and second kind. The period $L$ obtained from (25) and (26) reads
$L(\psi)=\frac{2\pi(G+\iota I)}{\iota_{N}}\left(B_{X}+(B_{M}-B_{X})\frac{E(m)}{K(m)}\right)^{-1}.$ (27)
The quantities $L$ and $\mathcal{D}$ are related through the periodicity of the $\mathrm{dn}$ function by
$\sqrt{\mathcal{D}}\,L(\psi)=\frac{4K(m)}{\sqrt{B_{M}-B_{X}}}.$ (28)
Substituting Eqs. (26) and (28) into the geometric constraint (25) and solving for $\sqrt{\mathcal{D}}$, we obtain
$\sqrt{\mathcal{D}}=\frac{\iota_{N}}{G+\iota I}\left(\frac{K}{\pi/2}\right)\frac{B_{X}+(B_{M}-B_{X})\,E/K}{\sqrt{B_{M}-B_{X}}}.$ (29)
Every factor on the right-hand side of Eq. (29) is either a flux function ($G$, II, $\iota$) or depends on $\psi$ only through the three roots, since $m$, $K(m)$, and $E(m)$ are all determined by $\{B_{M},B_{m},B_{X}\}$ alone. Therefore
$\mathcal{D}=\mathcal{D}\!\bigl(B_{M}(\psi),\,B_{m}(\psi),\,B_{X}(\psi),\,\iota(\psi)\bigr),$ (30)
and the description of quasisymmetric $B$ on a flux surface requires exactly three spectral functions of $\psi$ (the three roots of the cubic) plus the rotational transform profile, which is constrained independently by force balance. This is the precise statement of the hidden lower dimensionality of $B$.
We now show that the cubic formula for $(\partial_{\ell}B)^{2}$ given in (22) is connected to the KdV equation,
$\partial_{t}B+6B\partial_{x}B+\partial_{x}^{3}B=0.$ (31)
We differentiate (22) twice with respect to $\ell$ and use the $\partial_{\alpha}B$ equation (12) to obtain the KdV equation for $B$. We find that the coordinates $(x,t)$ are related to $(\alpha,\ell)$ through
$\frac{x}{\ell}=\sqrt{\frac{\mathcal{D}}{2}},\;\;\frac{t}{\int d\alpha H}=\frac{\sqrt{\mathcal{D}/2}}{2(B_{M}+B_{m}+B_{X})}.$ (32)
The $\partial_{\alpha}B$ relation (12), now takes the following form of a 1D TW equation with speed $c(\psi)$,
@@
Since the speed $c(\psi)$ is constant on a flux surface, we reiterate that quasisymmetric $B$ satisfying (31) must be periodic TW solutions of the KdV equation. We show this explicitly by rewriting the exact solution for $B$ (23) in terms of the cnoidal solution (Kamchatnov, 2000) of KdV,
$B$ $=B_{m}+(B_{M}-B_{m})\text{cn}^{2}\left(\sqrt{\frac{B_{M}-B_{X}}{2}}(x-ct),m\right).$ (34)
Alternatively, we can start with the KdV equation, move to the TW frame, and integrate the KdV equation twice to recover (22). The adiabatic invariant and the integrals (17) obtained from KdV potentials are manifestly time ($\alpha$) independent (Gjaja and Bhattacharjee, 1992; Berry and Howls, 1990; Brezhnev, 2008) because KdV shares the Painlevé property.
The Jacobi cosine elliptic function $\text{cn}(u,m)$ with modulus $m$ has maximum amplitude of unity. Thus, (34) confirms that $B$ varies between the minimum and the maximum values $B_{m},B_{M}$ of $B$. Two limiting cases, $m\to 0$ and $m\to 1$, are of particular interest. In the $m\to 0$ limit, $\mathrm{cn}(u,m)\to\cos(u)$ whereas in the $m\to 1$ limit $\mathrm{cn}(u,m)\to\mathrm{sech}(u)$. Therefore, $B$ is sinusoidal in the first limit and the KdV reflectionless potential (Gjaja and Bhattacharjee, 1992) in the second limit.
In the $m\to 0$ limit, both $B_{m}$ and $B_{M}$ approach a common value (say $B_{0}$) from (23) and $B$ is sinusoidal. This is the NAE limit, where, the on the axis $B$ is a constant, $B_{0}$, and varies as $B\approx B_{0}(1+\overline{\eta}\sqrt{2\psi}\cos\vartheta_{B})+2\psi(B_{20}+B_{22}\cos 2\vartheta_{B})$. The constant $\overline{\eta}$ dominates the maximum variation of $B$ on a flux surface, with the $O(2\psi)$ term being only a small second-order correction. Thus, the well-understood behavior of $B$ in Boozer coordinates near the magnetic axis (A. and H., 1991b, a; Landreman and Sengupta, 2018) is consistent with (22). As $m\to 0$, $E(m),K(m)$ both approach $\pi/2$, which implies that $L\to 2\pi(G+\iota I)/(B_{0}\iota_{N})$ and $\sqrt{D}\to\iota_{N}B_{0}/((G+\iota I)\sqrt{B_{0}-B_{X}})$ are both finite provided $B_{0}-B_{X}>0$ and $\iota_{N}\neq 0$. Further, the NAE predictions are
@@
Since $B_{22},B_{0},\overline{\eta}$ are all constants related to the $B$ spectrum in Boozer coordinates, within the NAE framework, $B_{X}$ and $\mathcal{D}$ are constant throughout the volume.
We now proceed to the $m\to 1$ limit. We caution that this limit must be interpreted with care. In this limit, $E\to 1$, but $K(m)\to\ln{4/\sqrt{1-m}}$ diverges logarithmically, and $B$ approaches a soliton profile. While the cnoidal solution (23) describes periodic functions with finite connection length, the soliton profile is non-periodic and corresponds to an infinite domain. Thus, this limit represents a breakdown of the periodicity assumption underlying QS in toroidal geometry. This has important consequences. Expressions for $L$ and $\sqrt{D}$ in (27) and (29) were obtained under the assumption that $B$ remains periodic with finite $L$, and therefore cannot be naively extrapolated to the $m\to 1$ limit, where $\sqrt{D}L$ diverges. The apparent finiteness of $L$ obtained by direct substitution reflects the use of formulae outside their domain of validity. Physically, the degeneration to a solitary-wave profile signals the loss of global periodicity and the breakdown of Boozer straight-field-line coordinates. In this regime, the connection length should be regarded as diverging or ceasing to be well-defined in the periodic sense. Since the non-periodic $m\to 1$ limit may not be correctly captured in our calculations, it is instructive to think about physical situations where periodicity is broken in stellarators.
Firstly, the limit in which $\iota_{N}\to 0$ corresponds to standard X-points in tokamaks, which are well understood. In this case, $L_{c}\sim 1/\iota$, and periodicity along a field line is lost because the field line cannot cross the X-points. This scaling is in fact correctly captured by (27), where, if $\iota_{N}$ or $B_{X}$ go to zero, $L$ will diverge while $\sqrt{D}\to 0$, since the divergence in $\sqrt{D}$ only logarithmic. If QS is exact, X-points can form only when $\iota$ resonates with the helicity of QS (Rodríguez and Bhattacharjee, 2021). However, even for minor QS errors, small resonant magnetic fields can lead to the formation of X-points and islands.
Secondly, for irrational $\iota$, sharp structures on the flux surfaces (”ridges”) can also arise (Bader et al., 2019; Punjabi and Boozer, 2020). These sharp structures can also hinder magnetic field lines from crossing them (Boozer, 2015). In both these cases, if the field lines can not cross the X-points or ridges then they can not physically connect two consecutive maxima or minima of $B$, leading to a divergence of the connection length.
The above physically motivated divergence of the connection length shows how the assumptions used to derive the expressions for $L,\sqrt{D}$ are no longer valid in the $m\to 1$ limit. Indeed, equation (25) presupposes that $L(\psi)$ is finite, since $B$ has a positive minimum, and the integral of such a function over an infinite domain cannot be finite. Therefore, the loss of periodicity in $B$ corresponds to a breakdown of global Boozer straight-field-line coordinates.
@@
Thus, while not mathematically proven within the context of our calculations, we speculate that the $m\to 1$ limit corresponds to the formation of an X-point or a sharp ridge on the flux surface. In Section V, we indeed observe the formation of an X-point as we approach the limiting $m\to 1$ flux surface. In practical terms, this means that QS stellarators may naturally form sharp edges as the volume of quasisymmetry is pushed to its limit, e.g., for compact devices. Sharp edges are essential for non-resonant divertors (Boozer, 2015; Punjabi and Boozer, 2020). This could potentially have significant beneficial consequences for stellarator optimization, since these divertors would be obtained naturally by optimizing for quasisymmetry in compact devices.
Interestingly, a similar singular limit also appears in the isodynamic limit, where the magnetic field curvature has cnoidal solutions (Schief, 2003). The $m=1$ limit is never achieved because of a geometrical obstruction when the central toroidal hole pinches. A similar pinching effect, leading to formation of cusps and break down of nestedness in quasiaxisymmetric devices, has been discussed in (Plunk, 2020; Brown et al., 2025). Other topological obstructions such as the “Reeb components” (Duignan et al., 2024) may arise before the $m\to 1$ limit is obtained. The Reeb case requires the toroidal magnetic field to vanish at some point if MHS force balance is imposed (Duignan et al., 2024).
V Numerical verification